The sum of two numbers is 240.The larger number is 6 less than twice the smaller. find the numbers
Answer by Paul(988)

Let the smaller number be x
"The larger number is 6 less that twice the smaller"= 2x-6
Sum = 240
Set the equation = 2x-6=240
=2x-6+6=240+6
=2x=246
=2x/2=246/2
x=123
the other number is 240-123 = 117
Therefore the numbers are 123 and 117